OpenStudy (anonymous):

at best buy, larry buys three cds and three dvds for $96 from the sale rack. Kelly buys two cds and one dvd from the same rack for $44. what is the price per cd and per dvd

OpenStudy (anonymous):

cds = c dvds = d 3c+3d = 96 2c+d=44 d = 44-2c sub it in 3c + 3(44-2c) = 96 3c+132-6c = 96 -3c = -36 c = 12 $12 per cd solve now for dvd d=44-2(12) d = 44-24 d = 20 $20 per dvd